OpenStudy (anonymous):

simplify each radical expression. use absolute value symbols when needed. show your work 1. square root of 25ysquared 2. square root of 49 x power of 4 3. cubic root square root of -8x power of 9 4. cubic root -0.125y power of 6

jimthompson5910 (jim_thompson5910):

I'll do the first one to get you started # 1 \[\Large \sqrt{25y^2} = \sqrt{25}*\sqrt{y^2}\] \[\Large \sqrt{25y^2} = \sqrt{5^2}*\sqrt{y^2}\] \[\Large \sqrt{25y^2} = 5|y|\]